# Large-Scale Model Enabled Semantic Communication Based on Robust Knowledge Distillation

### v1

@@ I-B Contributions
Simulation results show that the proposed RKD-SC framework preserves 95.86% of the performance of the teacher model while reducing the number of parameters by approximately 94.06% at an SNR of 25 dB and achieves performance gains exceeding 83.12% compared to the teacher model at an SNR of --10 dB on CIFAR10 dataset. Our results also demonstrate that the RKD-SC framework effectively transfers capabilities from large-scale to small-scale models while simultaneously maintaining robustness.
@@ III-A KD-based Lightweight DARTS
[…]
@@
[…]
@@
1:Teacher network $f_{\boldsymbol{\theta}_{\text{t}}}$, Training data $\mathcal{D}_{\text{train}}$, Validation data $\mathcal{D}_{\text{val}}$, Candidate operations $\mathcal{O}^{(l)}$ for layers $l=1...L$, Regularization $\lambda_{\mathcal{J}}>0$, Learning rates $\eta_{\alpha}$, $\eta_{\theta}$, Step size $\xi$, Number of operations $k$
@@
[…]
@@ IV-A1 Architecture and Hyperparameter
Input 224×224×3224$\times$ 224$\times$ 3224 × 224 × 3 - - - -
Stem Conv1 224×224×3224$\times$ 224$\times$ 3224 × 224 × 3 32 2 3x3 Conv, BN, ReLU Fixed
Stem Conv2 112×112×32112$\times$ 112$\times$ 32112 × 112 × 32 32 1 3x3 Conv, BN, ReLU Fixed
Stem Conv3 112×112×32112$\times$ 112$\times$ 32112 × 112 × 32 64 1 3x3 Conv, BN, ReLU Fixed
Stem Pool 112×112×64112$\times$ 112$\times$ 64112 × 112 × 64 64 2 2x2 AvgPool Fixed
Layer 1 56×56×6456$\times$ 56$\times$ 6456 × 56 × 64 64(16×4)64~(16$\times$ 4)64 ( 16 × 4 ) 1 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 2 56×56×12856$\times$ 56$\times$ 12856 × 56 × 128 128(32×4)128~(32$\times$ 4)128 ( 32 × 4 ) 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 3 28×28×25628$\times$ 28$\times$ 25628 × 28 × 256 256(64×4)256~(64$\times$ 4)256 ( 64 × 4 ) 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 4 14×14×51214$\times$ 14$\times$ 51214 × 14 × 512 512(128×4)512~(128$\times$ 4)512 ( 128 × 4 ) 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Head 7×7×10247$\times$ 7$\times$ 10247 × 7 × 1024 512512512 - AttentionPool2d Fixed (8 heads)
@@ IV-B1 Validation of RKD-SC
Fig. 6 presents the results of the RKD-SC framework, which leverages RKD algorithm to fine-tune the architecture identified by KDL-DARTS, compared with several baseline methods. As shown in Fig. 6, under AWGN channels at an SNR of 25 dB, the RKD-SC framework preserves 95.86%, 90.20%, and 78.39% of the performance of the teacher model (labeled as “JSCC-teacher” in Fig. 6) on the CIFAR10, CIFAR100, and ImageNet datasets, respectively, while significantly reducing the number of parameters by approximately 94.06%, 93.27%, and 93.26%. Moreover, RKD-SC outperforms the JSCC-student model by over 3.41%, 5.10%, and 12.25% on the three respective datasets. These results indicate that the RKD-SC framework effectively transfers the semantic representation capabilities from the teacher to the compact semantic encoder. Additionally, at an SNR of --10 dB, RKD-SC achieves performance gains exceeding 83.12%, 516.16%, and 107.51% compared to the JSCC-teacher on CIFAR10, CIFAR100, and ImageNet datasets, respectively. This demonstrates that the proposed CAT module significantly enhances the robustness against channel noise through the second stage of RKD algorithm within the RKD-SC framework.
@@ IV-B4 Ablation Study
Fig. 8 illustrates the impact of varying compression ratios within the CAT module. At an SNR of 10-10- 10 dB, for the CIFAR10 dataset, the system’s top-1 accuracy significantly improves from 11.73% to 63.80% as the compression ratio increases. Conversely, for the CIFAR100 and ImageNet datasets, the top-1 accuracy declines from 47.32% to 16.15% and from 12.76% to 7.47%, respectively, with an increasing compression ratio. This is because, in CAT, a higher compression ratio corresponds to less preservation of source information and a greater incorporation of channel features. For simpler datasets like CIFAR10, fewer semantic features are sufficient to represent the source information; thus, a higher compression ratio effectively enhances task performance under low SNR conditions by introducing rich channel features. However, for more complex datasets like CIFAR100 and ImageNet, richer semantic representations are essential. Although higher compression ratios introduce additional channel features, the substantial loss of critical source information negatively impacts overall task performance.

### v2

@@ I-B Contributions
Simulation results show that the proposed RKD-SC framework preserves 95.86% of the performance of the teacher model while reducing the number of parameters by approximately 94.06% at an SNR of 25 dB and achieves performance gains exceeding 83.12% compared to the teacher model at an SNR of -10 dB on CIFAR10 dataset. Our results also demonstrate that the RKD-SC framework effectively transfers capabilities from large-scale to small-scale models while simultaneously maintaining robustness.
@@ III-A KD-based Lightweight DARTS
Figure 3: An overview of the proposed KDL-DARTS method: (a) The operations between layer $l$ and layer $l+1$ are initially undetermined. (b) Candidate operations are introduced, and their optimal continuous operation weights, denoted as $\boldsymbol{\alpha}^{(l)}$, are derived by solving a bilevel optimization problem. (c) A combined metric is computed using the Hadamard product: $\boldsymbol{\alpha}^{(l)}\circ(1-\boldsymbol{\beta}^{(l)})$. (d) The top-$k$ optimal operations are selected based on the values obtained from this combined metric.
@@
Consistent with the DARTS framework, we introduce a vector of continuous candidate operation weights $\boldsymbol{\alpha}^{(l)}=\{\alpha^{(l,j)}\}\subsetneq\boldsymbol{\alpha}$ to combine outputs from all candidate operations between the $l$-th and the $(l+1)$-th layers:
@@
1:Teacher network $f_{\boldsymbol{\theta}_{\text{t}}}$, Training data $\mathcal{D}_{\text{train}}$, Validation data $\mathcal{D}_{\text{val}}$, Candidate operations $\mathcal{O}^{(l)}$ for layers l=1Ll=1...L, Regularization $\lambda_{\mathcal{J}}>0$, Learning rates $\eta_{\alpha}$, $\eta_{\theta}$, Step size $\xi$, Number of operations $k$
@@
Upon completion of the training process, we obtain the optimal architecture parameters $\boldsymbol{\alpha}^{*}$ and the corresponding model parameters $\boldsymbol{\theta}^{*}$. The final selection of candidate operations is determined jointly by $\boldsymbol{\alpha}^{*}$ and the penalty factors $\boldsymbol{\beta}^{*}$. Specifically, we select the top-$k$ operations based on the combined metric $\alpha^{(l,j)}\cdot(1-\beta^{(l,j)})$, as illustrated in Fig. 3, thereby deriving the optimal architecture $a^{*}$. Here, the term $(1-\beta^{(l,j)})$ is adopted instead of $\beta^{(l,j)}$ to effectively penalize candidate operations with larger parameter quantities. Unlike the original DARTS approach, where the final model parameters $\boldsymbol{\theta}^{*}$ are directly employed following the optimization process, our proposed KDL-DARTS framework solely yields the optimal architecture $a_{i}^{*}$ of the $i$-th semantic encoder. In the following sections, we explore how the semantic representation capability of a large-scale model can be effectively transferred to the semantic encoder while maintaining robustness.
@@ IV-A1 Architecture and Hyperparameter
Input $224\times 224\times 3$ - - - -
Stem Conv1 $224\times 224\times 3$ 32 2 3x3 Conv, BN, ReLU Fixed
Stem Conv2 $112\times 112\times 32$ 32 1 3x3 Conv, BN, ReLU Fixed
Stem Conv3 $112\times 112\times 32$ 64 1 3x3 Conv, BN, ReLU Fixed
Stem Pool $112\times 112\times 64$ 64 2 2x2 AvgPool Fixed
Layer 1 $56\times 56\times 64$ $64~(16\times 4)$ 1 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 2 $56\times 56\times 128$ $128~(32\times 4)$ 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 3 $28\times 28\times 256$ $256~(64\times 4)$ 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Layer 4 $14\times 14\times 512$ $512~(128\times 4)$ 2 MixedLayer(Bottleneck) Depth Search: $k\in\{1..5\}$ blocks
Head $7\times 7\times 1024$ 512512 - AttentionPool2d Fixed (8 heads)
@@ IV-B1 Validation of RKD-SC
Fig. 6 presents the results of the RKD-SC framework, which leverages RKD algorithm to fine-tune the architecture identified by KDL-DARTS, compared with several baseline methods. As shown in Fig. 6, under AWGN channels at an SNR of 25 dB, the RKD-SC framework preserves 95.86%, 90.20%, and 78.39% of the performance of the teacher model (labeled as “JSCC-teacher” in Fig. 6) on the CIFAR10, CIFAR100, and ImageNet datasets, respectively, while significantly reducing the number of parameters by approximately 94.06%, 93.27%, and 93.26%. Moreover, RKD-SC outperforms the JSCC-student model by over 3.41%, 5.10%, and 12.25% on the three respective datasets. These results indicate that the RKD-SC framework effectively transfers the semantic representation capabilities from the teacher to the compact semantic encoder. Additionally, at an SNR of -10 dB, RKD-SC achieves performance gains exceeding 83.12%, 516.16%, and 107.51% compared to the JSCC-teacher on CIFAR10, CIFAR100, and ImageNet datasets, respectively. This demonstrates that the proposed CAT module significantly enhances the robustness against channel noise through the second stage of RKD algorithm within the RKD-SC framework.
@@ IV-B4 Ablation Study
Fig. 8 illustrates the impact of varying compression ratios within the CAT module. At an SNR of 10-10 dB, for the CIFAR10 dataset, the system’s top-1 accuracy significantly improves from 11.73% to 63.80% as the compression ratio increases. Conversely, for the CIFAR100 and ImageNet datasets, the top-1 accuracy declines from 47.32% to 16.15% and from 12.76% to 7.47%, respectively, with an increasing compression ratio. This is because, in CAT, a higher compression ratio corresponds to less preservation of source information and a greater incorporation of channel features. For simpler datasets like CIFAR10, fewer semantic features are sufficient to represent the source information; thus, a higher compression ratio effectively enhances task performance under low SNR conditions by introducing rich channel features. However, for more complex datasets like CIFAR100 and ImageNet, richer semantic representations are essential. Although higher compression ratios introduce additional channel features, the substantial loss of critical source information negatively impacts overall task performance.